OpenStudy (anonymous):

Find the difference quotient (a). f(x) - f(a) / x-a and for (b). f(x+h) - f(x) / h if f(x) = 2x^2 - x +1

OpenStudy (anonymous):

need help on b) ?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

plug x+h into f(x) everywhere there's an 'x'

OpenStudy (anonymous):

eg.: f(x) = x^4 +2x f(x+h) = (x+h)^4 +2(x+h)

OpenStudy (anonymous):

so 2(x+h)^2 - x(x+h) + 1 - (2x^2 - x + 1) / h ?

OpenStudy (anonymous):

2(x+h)^2 - x(x+h) + 1 - (2x^2 - x + 1) / h ? ^-----^ ...?

OpenStudy (anonymous):

you have an extra 'x' there...

OpenStudy (anonymous):

ahh so when it is just x.. u dont have to reuse the whole term inside the parens?

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

if it's 'x', simply sub.s in "x+h"

OpenStudy (anonymous):

so does it end up being 2(x+h)(x+h)-x-h-2x^2-x?

OpenStudy (anonymous):

then expand and simplify.... there will be some terms that cancel...

OpenStudy (anonymous):

and you never end up redistributing that 2?

OpenStudy (anonymous):

Expand (x+h)(x+h)

OpenStudy (anonymous):

2(x^2+2xh+h^2) - x - h ....

OpenStudy (anonymous):

2(x^2+2xh+h^2) -x-h-2x^2+x

OpenStudy (anonymous):

2x^2+4xh+2h^2 -x -h -2x^2 +x

OpenStudy (anonymous):

lol im gonna run out of eraser :)

OpenStudy (anonymous):

4x-2h? final answer?

OpenStudy (anonymous):

2h^2 +4xh -h

OpenStudy (anonymous):

\[ \frac{ 2h^2 +4xh -h }{ h }\]

OpenStudy (anonymous):

h in every term...

OpenStudy (anonymous):

so pull out the h?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

only in the numerator and divide the demonator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok!!! awesome thanks so much!!!

OpenStudy (anonymous):

sure:)